# Distributed Cooperative AI for Large-Scale Eigenvalue Computations Using Neural Networks

### v1

@@ Abstract
This paper presents a novel method for eigenvalue computation using a distributed cooperative neural network framework. Unlike traditional techniques that struggle with scalability in large systems, our decentralized algorithm enables multiple autonomous agents to collaboratively estimate the smallest eigenvalue of large matrices. Each agent uses a localized neural network model, refining its estimates through inter-agent communication. Our approach guarantees convergence to the true eigenvalue, even with communication failures or network disruptions. Theoretical analysis confirms the robustness and accuracy of the method, while empirical results demonstrate its better performance compared to some traditional centralized algorithms
Keywords: Distributed Neural Networks, Eigenvalue Computation, Decentralized Algorithms, Cooperative AI, Scalability and Robustness
@@ 2 Preliminaries
2.1 Eigenvalue Problems
For a square matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, an eigenvalue $\lambda\in\mathbb{C}$ and its corresponding eigenvector $\mathbf{v}\in\mathbb{C}^{n}$ satisfy the equation
$\mathbf{A}\mathbf{v}=\lambda\mathbf{v},\quad\mathbf{v}\neq\mathbf{0}.$
Computing eigenvalues, especially for large matrices, is computationally demanding. Traditional methods like the QR algorithm and power iteration are typically centralized, requiring significant memory and processing resources. In distributed systems, the matrix $\mathbf{A}$ can be partitioned into submatrices $\mathbf{A}_{i}$, such that
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}&\mathbf{A}_{2}&\dots&\mathbf{A}_{m}\end{bmatrix}.$
Each agent $i$ operates on its submatrix $\mathbf{A}_{i}$ and, through iterative communication with other agents, collaboratively approximates the eigenvalues of $\mathbf{A}$.
@@ 2.2 Neural Networks for Eigenvalue Approximation
Neural networks (NNs) are increasingly used to approximate complex functions that are computationally intensive. Consider a neural network $f_{\theta}(\mathbf{A})$ parameterized by $\theta$, where
$f_{\theta}(\mathbf{A})\approx\lambda(\mathbf{A}).$
Training the network involves minimizing a loss function $\mathcal{L}$ that quantifies the difference between the predicted and actual eigenvalues
$\mathcal{L}(\theta)=\sum_{i=1}^{N}\left|f_{\theta}(\mathbf{A}_{i})-\lambda_{i}\right|^{2},$
where $\lambda_{i}$ are the true eigenvalues for a set of training matrices $\{\mathbf{A}_{i}\}$. In the proposed distributed framework, each agent uses its NN to approximate eigenvalues based on its local submatrix $\mathbf{A}_{i}$. Agents iteratively share intermediate results to refine their estimates. The update rule is given by
$\lambda_{i}^{(k+1)}=f_{\theta_{i}}(\mathbf{A}_{i})+\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}(\lambda_{j}^{(k)}-\lambda_{i}^{(k)}),$
where $\mathcal{N}_{i}$ denotes the neighboring agents, $\alpha_{ij}$ are weighting factors, and $k$ is the iteration index. This approach enables the system to converge to the global eigenvalues through local interactions. The key challenges include ensuring algorithmic convergence, maintaining accuracy comparable to centralized methods, and optimizing communication overhead. This research aims to develop a scalable and efficient distributed algorithm that integrates neural networks to address large-scale eigenvalue problems effectively.
@@ 3 Methodology
In this section, we describe the methodology for implementing distributed cooperative AI for large-scale eigenvalue computations using neural networks. The methodology involves three primary components
  1. (a)
    Neural Network Design: A neural network architecture designed to approximate eigenvalues from submatrices.
  2. (b)
    Distributed Framework: A cooperative framework where multiple agents work together, sharing information to converge on the global eigenvalues.
  3. (c)
    Algorithm Implementation: The integration of neural network predictions with the distributed system to iteratively refine the eigenvalue estimates.
@@ 3.1 Neural Network Design
The neural network $f_{\theta}$ is designed to approximate the eigenvalue function $\lambda(\mathbf{A})$ for any given square matrix $\mathbf{A}$. The architecture is chosen to balance complexity and generalization ability, typically using fully connected layers with activation functions that handle nonlinearity. The network is trained on a dataset of matrices $\mathbf{A}_{i}$ and their corresponding eigenvalues $\lambda_{i}$, minimizing the loss function
$\mathcal{L}(\theta)=\sum_{i=1}^{N}\left\|f_{\theta}(\mathbf{A}_{i})-\lambda_{i}\right\|^{2},$
where $N$ is the number of training samples. The training process involves standard backpropagation and stochastic gradient descent (SGD) or its variants.
@@ 3.2 Distributed Framework
The distributed framework involves multiple agents, each responsible for a subset of the global matrix $\mathbf{A}$. The matrix $\mathbf{A}$ is partitioned into $m$ submatrices $\mathbf{A}_{i}$ such that
$\mathbf{A}=\bigoplus_{i=1}^{m}\mathbf{A}_{i}.$
Each agent $i$ maintains a local copy of $\mathbf{A}_{i}$ and an associated neural network $f_{\theta_{i}}$. The agents communicate iteratively, sharing their local eigenvalue estimates to refine their predictions collectively. The communication model follows a cooperative scheme, where each agent updates its estimate based on its neighbors’ estimates.
@@ 3.3 Algorithm Implementation
The following algorithm outlines the distributed cooperative eigenvalue computation using neural networks.
Algorithm 1 Distributed Cooperative Eigenvalue Computation Using Neural Networks
0:  Matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, number of agents $m$, neural networks $f_{\theta_{1}},f_{\theta_{2}},\dots,f_{\theta_{m}}$, maximum iterations $K$, convergence tolerance $\epsilon$
0:  Eigenvalue estimates $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$
1:  Partition $\mathbf{A}$ into submatrices $\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{m}$
2:  Initialize local eigenvalue estimates $\lambda_{i}^{(0)}=f_{\theta_{i}}(\mathbf{A}_{i})$ for all agents $i=1,\dots,m$
3:  Set iteration counter $k=0$
4:  repeat
5:     for each agent $i=1,\dots,m$ do
6:        Receive eigenvalue estimates $\lambda_{j}^{(k)}$ from neighboring agents $j\in\mathcal{N}_{i}$
7:        Update local estimate
$\lambda_{i}^{(k+1)}=f_{\theta_{i}}(\mathbf{A}_{i})+\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}(\lambda_{j}^{(k)}-\lambda_{i}^{(k)})$
8:     end for
9:     Increment iteration counter $k=k+1$
10:  until convergence $\|\lambda_{i}^{(k+1)}-\lambda_{i}^{(k)}\|<\epsilon$ for all agents $i$ or $k\geq K$
11:  Aggregate final eigenvalue estimates from all agents $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$
The matrix $\mathbf{A}$ is partitioned into submatrices, and each agent initializes its local eigenvalue estimate using its neural network. Then, the agents enter an iterative loop where they share their eigenvalue estimates with their neighbors and update their local estimates based on the received information. The loop continues until the estimates converge or the maximum number of iterations is reached. Finally, the final eigenvalue estimates are aggregated from all agents.
3.4 Convergence and Scalability
The convergence of the algorithm is influenced by the choice of the weighting factors $\alpha_{ij}$ and the structure of the communication graph $\mathcal{N}_{i}$. The method is scalable, as it can handle larger matrices by increasing the number of agents, each with its local neural network, ensuring that the computational load is distributed across the system.
@@ 4 Theoretical Results
In this section, we provide a comprehensive theoretical analysis of the proposed distributed cooperative eigenvalue computation method using neural networks. The analysis includes proofs of convergence, error bounds, and discussions on computational complexity and communication overhead.
@@ 4.1 Convergence Analysis
We aim to demonstrate that the iterative algorithm converges to the true eigenvalues of the matrix $\mathbf{A}$ under certain conditions. Now, we make the following assumptions
  1. (i)
    the global matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ is symmetric and positive definite.
  2. (ii)
    the communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ among agents is connected.
  3. (iii)
    the weighting factors $\alpha_{ij}$ satisfy
    $\alpha_{ij}=\alpha_{ji}\geq 0,\quad\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}=1$
  4. (iv)
    the neural networks $f_{\theta_{i}}$ provide unbiased estimators of the local eigenvalues with bounded variance
    $\mathbb{E}[f_{\theta_{i}}(\mathbf{A}_{i})]=\lambda_{i}^{*},\quad\text{Var}[f_{\theta_{i}}(\mathbf{A}_{i})]\leq\sigma^{2}$
@@
The consensus error at iteration $k$ is defined as $e^{(k)}=\max_{i,j}\|\lambda_{i}^{(k)}-\lambda_{j}^{(k)}\|$.
@@
Under assumptions (i) - (iii), the sequence $\{\lambda_{i}^{(k)}\}$ generated by the algorithm converges exponentially to a consensus value $\lambda^{(k)}$ as $k\to\infty$
$e^{(k)}\leq\rho^{k}e^{(0)},\quad\rho=1-\delta$
, where $10<\delta<$ 1 depends on the spectral properties of the Laplacian matrix of $\mathcal{G}$.
@@
Consider the update rule
$\lambda_{i}^{(k+1)}=f_{\theta_{i}}(\mathbf{A}_{i})+\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}(\lambda_{j}^{(k)}-\lambda_{i}^{(k)}).$
Define the aggregate vector $\boldsymbol{\lambda}^{(k)}=[\lambda_{1}^{(k)},\lambda_{2}^{(k)},\dots,\lambda_{m}^{(k)}]^{T}$. The update can be written in matrix form
$\boldsymbol{\lambda}^{(k+1)}=\mathbf{W}\boldsymbol{\lambda}^{(k)}+\mathbf{f},$
where $\mathbf{W}$ is the weight matrix derived from $\alpha_{ij}$, and $\mathbf{f}=[f_{\theta_{1}}(\mathbf{A}_{1}),\dots,f_{\theta_{m}}(\mathbf{A}_{m})]^{T}$. Since $\mathcal{G}$ is connected and $\alpha_{ij}$ satisfy the given conditions, $\mathbf{W}$ is doubly stochastic and its second largest eigenvalue modulus (SLEM) $\rho$ satisfies $10\leq\rho<$ 1. Applying iterative updates
$\boldsymbol{\lambda}^{(k)}=\mathbf{W}^{k}\boldsymbol{\lambda}^{(0)}+\sum_{t=0}^{k-1}\mathbf{W}^{t}\mathbf{f}.$
The consensus error becomes
$e^{(k)}=\|\boldsymbol{\lambda}^{(k)}-\bar{\lambda}^{(k)}\mathbf{1}\|\leq\rho^{k}e^{(0)}+C,$
where $\bar{\lambda}^{(k)}$ is the average estimate and $C$ is bounded due to the bounded variance of $f_{\theta_{i}}$. As $k\to\infty$, $e^{(k)}\to 0$, proving convergence to consensus.∎
@@
If the neural network estimators are consistent, i.e.,
$\lim_{N\to\infty}f_{\theta_{i}}(\mathbf{A}_{i})=\lambda_{i}^{*}$
then the consensus value converges to the true eigenvalues
$\lim_{k\to\infty}\lambda_{i}^{(k)}=\lambda^{*},\quad\forall i$
The convergence rate depends on the spectral gap of $\mathbf{W}$. A larger spectral gap implies faster convergence. The choice of $\alpha_{ij}$ can be optimized to maximize this gap.

### v2

# Decentralized Neural Networks for Robust and Scalable Eigenvalue Computation

@@ Abstract
This paper introduces a novel method for eigenvalue computation using a distributed cooperative neural network framework. Unlike traditional techniques that face scalability challenges in large systems, our decentralized algorithm enables multiple autonomous agents to collaboratively estimate the smallest eigenvalue of large matrices. Each agent employs a localized neural network, refining its estimates through communication with neighboring agents. Our empirical results confirm the algorithm’s convergence towards the true eigenvalue, with estimates clustered closely around the true value. Even in the presence of communication delays or network disruptions, the method demonstrates strong robustness and scalability. Theoretical analysis further validates the accuracy and stability of the proposed approach, while empirical tests highlight its efficiency and precision, surpassing traditional centralized algorithms in large-scale eigenvalue computations.
Keywords: Distributed Neural Networks, Eigenvalue Computation, Decentralized Algorithms, Cooperative AI, Scalability, Robustness
@@ 2 Preliminaries
2.1 Eigenvalue Problems and Matrix Partitioning
Given a square matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$, the eigenvalue problem is to find scalar values $\lambda\in\mathbb{C}$ and non-zero vectors $\mathbf{v}\in\mathbb{C}^{n}$ such that
$\mathbf{A}\mathbf{v}=\lambda\mathbf{v}.$
For large-scale matrices, direct computation of eigenvalues can be computationally expensive. Therefore, in distributed systems, we partition $\mathbf{A}$ into smaller submatrices that can be processed by different agents in parallel.
2.1.1 Matrix Partitioning
Suppose we partition $\mathbf{A}$ into $m$ block submatrices $\mathbf{A}_{i}$ along its rows or columns. For simplicity, consider a row-wise partitioning:
$\mathbf{A}=\begin{bmatrix}\mathbf{A}^{(1)}\\ \mathbf{A}^{(2)}\\ \vdots\\ \mathbf{A}^{(m)}\end{bmatrix},$
where each submatrix $\mathbf{A}^{(i)}\in\mathbb{R}^{k_{i}\times n}$ and $\sum_{i=1}^{m}k_{i}=n$. Here, each submatrix $\mathbf{A}^{(i)}$ has $k_{i}$ rows but retains the same number of columns $n$ as the original matrix $\mathbf{A}$.
2.1.2 Eigenvalues of Submatrices
Each submatrix $\mathbf{A}^{(i)}$ can have up to $k_{i}$ eigenvalues, denoted as $\{\lambda_{j}^{(i)}\}_{j=1}^{k_{i}}$. In the distributed framework, we aim to approximate the $n$ eigenvalues $\{\lambda_{1},\lambda_{2},\dots,\lambda_{n}\}$ of the original matrix $\mathbf{A}$ by aggregating the computations from the submatrices. While each submatrix $\mathbf{A}^{(i)}$ provides a partial view of $\mathbf{A}$, the goal is for the collective computation of all submatrices to approximate the global eigenvalues of $\mathbf{A}$.
@@ 2.2 Neural Networks for Eigenvalue Approximation
2.2.1 Training Neural Networks for Submatrices
Each submatrix $\mathbf{A}^{(i)}\in\mathbb{R}^{k_{i}\times n}$ is assigned to an agent $i$ that trains a neural network $f_{\theta_{i}}$ to approximate the local eigenvalues. The objective is to train this neural network such that
$f_{\theta_{i}}(\mathbf{A}^{(i)})\approx\{\lambda_{j}^{(i)}\}_{j=1}^{k_{i}},$
where $\lambda_{j}^{(i)}$ are the eigenvalues of the submatrix $\mathbf{A}^{(i)}$.
Loss Function for Neural Network Training
The neural network $f_{\theta_{i}}$ is trained by minimizing a loss function that measures the discrepancy between the predicted eigenvalues and the true eigenvalues of the submatrix $\mathbf{A}^{(i)}$
$\mathcal{L}(\theta_{i})=\sum_{j=1}^{k_{i}}\left|f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}-\lambda_{j}^{(i)}\right|^{2}.$
Here, $f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}$ denotes the $j$-th predicted eigenvalue of the submatrix $\mathbf{A}^{(i)}$.
Parameter Fine-tuning
The parameters $\theta_{i}$ of the neural network $f_{\theta_{i}}$ are updated using gradient descent:
$\theta_{i}^{(t+1)}=\theta_{i}^{(t)}-\eta\nabla_{\theta_{i}}\mathcal{L}(\theta_{i}^{(t)}),$
where $\eta$ is the learning rate, and $\nabla_{\theta_{i}}\mathcal{L}(\theta_{i}^{(t)})$ is the gradient of the loss function with respect to $\theta_{i}$ at iteration $t$. This process iteratively refines the parameters $\theta_{i}$ to improve the approximation of the eigenvalues $\{\lambda_{j}^{(i)}\}$ for the submatrix $\mathbf{A}^{(i)}$.
2.2.2 Collaborative Eigenvalue Estimation
Once each neural network $f_{\theta_{i}}$ is trained on its submatrix $\mathbf{A}^{(i)}$, the agents begin a collaborative process to approximate the global eigenvalues of the original matrix $\mathbf{A}$. Each agent $i$ estimates the eigenvalues $\{\lambda_{j}^{(i)}\}$ of its submatrix and shares these estimates with neighboring agents.
Update Rule for Submatrix Eigenvalues
Let $\lambda_{j}^{(i,k)}$ be the estimate of the $j$-th eigenvalue of the submatrix $\mathbf{A}^{(i)}$ at iteration $k$. The update rule for refining these submatrix eigenvalues using the neural network and neighboring agents’ information is
$\lambda_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}+\sum_{l\in\mathcal{N}_{i}}\alpha_{il}\left(\lambda_{j}^{(l,k)}-\lambda_{j}^{(i,k)}\right).$
Here
  • $f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}$ is the $j$-th eigenvalue prediction by the neural network for submatrix $\mathbf{A}^{(i)}$.
  • $\mathcal{N}_{i}$ denotes the neighboring agents communicating with agent $i$.
  • $\alpha_{il}$ are weighting factors that determine the influence of agent $l$’s eigenvalue estimates on agent $i$.
This update rule ensures that the eigenvalue estimates of each submatrix converge through local interactions toward the global eigenvalues of the original matrix $\mathbf{A}$.
Global Eigenvalue Update Rule
To approximate the eigenvalues $\{\lambda_{j}\}_{j=1}^{n}$ of the original matrix $\mathbf{A}$, we combine the refined estimates from all agents. Define the global eigenvalue estimate $\Lambda_{j}^{(k)}$ at iteration $k$ as
$\Lambda_{j}^{(k+1)}=\sum_{i=1}^{m}\beta_{i}\lambda_{j}^{(i,k+1)},$
where
  • $\lambda_{j}^{(i,k+1)}$ are the refined eigenvalue estimates from each submatrix.
  • $\beta_{i}$ are weighting factors that reflect the contribution of each submatrix $\mathbf{A}^{(i)}$ to the global estimate.
The goal is for the global eigenvalue estimates $\{\Lambda_{j}^{(k)}\}_{j=1}^{n}$ to converge to the actual eigenvalues of the original matrix $\mathbf{A}$ as $k\to\infty$.
@@ 3 Methodology
This section outlines the methodology for implementing distributed cooperative AI for large-scale eigenvalue computations using neural networks. The methodology consists of three main components
  1. 1.
    Neural Network Design: Designing neural networks to approximate eigenvalues from submatrices.
  2. 2.
    Distributed Framework: A cooperative framework where multiple agents collaborate to converge on the global eigenvalues.
  3. 3.
    Algorithm Implementation: Integrating neural network predictions into a distributed system to iteratively refine eigenvalue estimates.
@@ 3.1 Neural Network Design
Each neural network $f_{\theta_{i}}$ is designed to approximate the eigenvalues of a submatrix $\mathbf{A}^{(i)}$. The architecture, typically a fully connected network with non-linear activation functions, is trained to learn the mapping
$f_{\theta_{i}}(\mathbf{A}^{(i)})\approx\{\lambda_{j}^{(i)}\}_{j=1}^{k_{i}},$
where $\lambda_{j}^{(i)}$ represents the eigenvalues of the submatrix $\mathbf{A}^{(i)}$. The network is trained on a dataset of submatrices $\mathbf{A}^{(i)}$ and their corresponding eigenvalues, using the loss function
$\mathcal{L}(\theta_{i})=\sum_{j=1}^{k_{i}}\left\|f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}-\lambda_{j}^{(i)}\right\|^{2}.$
By minimizing this loss function, the network learns to approximate the eigenvalue spectrum of the submatrix.
@@ 3.2 Distributed Framework
In the distributed framework, multiple agents, each assigned to a submatrix $\mathbf{A}^{(i)}$, collaborate to approximate the global eigenvalues of the matrix $\mathbf{A}$. Each agent trains its neural network and updates its eigenvalue estimates through local communication with neighboring agents. The update rule for submatrix eigenvalue estimates is
$\lambda_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}+\sum_{l\in\mathcal{N}_{i}}\alpha_{il}\left(\lambda_{j}^{(l,k)}-\lambda_{j}^{(i,k)}\right),$
and the global eigenvalue estimate is updated as
$\Lambda_{j}^{(k+1)}=\sum_{i=1}^{m}\beta_{i}\lambda_{j}^{(i,k+1)}.$
@@ 3.3 Algorithm Implementation
The overall algorithm proceeds as follows;
1:Input: Matrix $\mathbf{A}$, number of agents $m$, initial neural network parameters $\{\theta_{i}\}_{i=1}^{m}$, learning rates $\alpha_{il}$, aggregation weights $\beta_{i}$
2:Output: Estimated global eigenvalues $\{\Lambda_{j}\}$
3:Partition the matrix $\mathbf{A}$ into submatrices $\mathbf{A}^{(i)}$ for each agent $i$, $i=1,\ldots,m$.
4:for each agent $i$ do
5:     Initialize the neural network $f_{\theta_{i}}$ for submatrix $\mathbf{A}^{(i)}$.
6:     Train $f_{\theta_{i}}$ on submatrix $\mathbf{A}^{(i)}$.
7:end for
8:repeat
9:     for each agent $i$ do
10:         for each eigenvalue $j$ do
11:              Update the submatrix eigenvalue estimate
$\lambda_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}+\sum_{l\in\mathcal{N}_{i}}\alpha_{il}\left(\lambda_{j}^{(l,k)}-\lambda_{j}^{(i,k)}\right)$
12:         end for
13:     end for
14:     Combine estimates from all agents to approximate the global eigenvalues
$\Lambda_{j}^{(k+1)}=\sum_{i=1}^{m}\beta_{i}\lambda_{j}^{(i,k+1)}$
15:until global eigenvalue estimates $\{\Lambda_{j}^{(k)}\}$ converge
@@ 4 Theoretical Results
This section provides a comprehensive theoretical analysis of the proposed distributed cooperative eigenvalue computation method using neural networks. The analysis includes proofs of convergence, error bounds, and discussions on computational complexity and communication overhead.
@@ 4.1 Convergence Analysis
In this section, we demonstrate that the iterative algorithm converges to the true eigenvalues of the matrix $\mathbf{A}$ under certain conditions. The following assumptions are made;
  1. 1.
    Matrix Properties: The global matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ is symmetric and positive definite, meaning all its eigenvalues are real and positive.
  2. 2.
    Communication Graph: The communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ among agents is connected. This ensures that information can flow between any two agents, either directly or through intermediary agents.
  3. 3.
    Weighting Factors: The weighting factors $\alpha_{il}$ used in communication between agents satisfy:
    $\alpha_{il}=\alpha_{li}\geq 0,\quad\sum_{l\in\mathcal{N}_{i}}\alpha_{il}=1,$
    where $\mathcal{N}_{i}$ is the set of neighboring agents for agent $i$. These conditions ensure that the weight matrix used for updating eigenvalue estimates is symmetric and properly normalized.
  4. 4.
    Neural Network Estimators: The neural networks $f_{\theta_{i}}$ employed by each agent provide unbiased estimates of the local eigenvalues, with a variance bounded by $\sigma^{2}$
    $\mathbb{E}[f_{\theta_{i}}(\mathbf{A}^{(i)})]=\lambda_{i}^{*},\quad\text{Var}[f_{\theta_{i}}(\mathbf{A}^{(i)})]\leq\sigma^{2}.$
@@
This measures the difference between eigenvalue estimates from different agents at iteration $k$
$e^{(k)}=\max_{i,j}\|\lambda_{j}^{(i,k)}-\lambda_{j}^{(l,k)}\|.$
A smaller consensus error means that agents’ estimates are closer to agreement.
@@
Under assumptions (i)–(iii) in section 4.1 above, the sequence of eigenvalue estimates $\{\lambda_{j}^{(i,k)}\}$ generated by the algorithm converges exponentially to a consensus value $\lambda_{j}^{(k)}$ as $k\to\infty$
$e^{(k)}\leq\rho^{k}e^{(0)},\quad\rho=1-\delta,$
where $10<\delta<$ 1 depends on the spectral properties of the Laplacian matrix of the communication graph $\mathcal{G}$.
@@
The update rule for eigenvalue estimates at agent $i$ and iteration $k+1$ is given by
$\lambda_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})_{j}+\sum_{l\in\mathcal{N}_{i}}\alpha_{il}\left(\lambda_{j}^{(l,k)}-\lambda_{j}^{(i,k)}\right).$
Let $\boldsymbol{\lambda}_{j}^{(k)}=[\lambda_{j}^{(1,k)},\lambda_{j}^{(2,k)},\dots,\lambda_{j}^{(m,k)}]^{T}$ represent the eigenvalue estimates across all agents. This update can be rewritten in matrix form as
$\boldsymbol{\lambda}_{j}^{(k+1)}=\mathbf{W}\boldsymbol{\lambda}_{j}^{(k)}+\mathbf{f}_{j},$
where $\mathbf{W}$ is the weight matrix derived from $\alpha_{il}$, and $\mathbf{f}_{j}$ is the vector of neural network outputs from all agents. Since the communication graph is connected and $\alpha_{il}$ satisfies the conditions, $\mathbf{W}$ is doubly stochastic. Its second largest eigenvalue modulus (SLEM) $\rho$ satisfies $10\leq\rho<$ 1. After applying iterative updates, the consensus error satisfies
$e^{(k)}=\|\boldsymbol{\lambda}_{j}^{(k)}-\bar{\lambda}_{j}^{(k)}\mathbf{1}\|\leq\rho^{k}e^{(0)}+C,$
where $\bar{\lambda}_{j}^{(k)}$ is the average eigenvalue estimate, and $C$ is bounded due to the neural network’s variance. As $k\to\infty$, $e^{(k)}\to 0$, hence the convergence to consensus.∎
@@
If the neural network estimators are consistent, i.e., as the training data increases
$\lim_{N\to\infty}f_{\theta_{i}}(\mathbf{A}^{(i)})=\lambda_{j}^{*},$
then the consensus value converges to the true eigenvalue
$\lim_{k\to\infty}\lambda_{j}^{(i,k)}=\lambda_{j}^{*},\quad\forall i.$
Proof.
Consider the update rule for the eigenvalue estimate $\lambda_{j}^{(i,k+1)}$ at time step $k+1$ for agent $i$
$\lambda_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})+\sum_{l\in\mathcal{N}_{i}^{(k)}}w_{il}^{(k)}(\lambda_{j}^{(l,k)}-\lambda_{j}^{(i,k)}),$
where $\mathcal{N}_{i}^{(k)}$ is the set of neighbors of agent $i$ in the communication graph at time $k$, and $w_{il}^{(k)}$ are the weighting coefficients satisfying $\sum_{l\in\mathcal{N}_{i}^{(k)}}w_{il}^{(k)}=1$. Define the error at time $k$ for agent $i$ as the difference between the current eigenvalue estimate $\lambda_{j}^{(i,k)}$ and the true eigenvalue $\lambda_{j}^{*}$:
$e_{j}^{(i,k)}=\lambda_{j}^{(i,k)}-\lambda_{j}^{*}.$
The update rule for the error can be written as
$e_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})-\lambda_{j}^{*}+\sum_{l\in\mathcal{N}_{i}^{(k)}}w_{il}^{(k)}(e_{j}^{(l,k)}-e_{j}^{(i,k)}).$
Since $f_{\theta_{i}}(\mathbf{A}^{(i)})\to\lambda_{j}^{*}$ as $N\to\infty$ (by the consistency of the neural network estimator), for sufficiently large $N$, we can approximate
$f_{\theta_{i}}(\mathbf{A}^{(i)})=\lambda_{j}^{*}+\epsilon_{i},$
where $\epsilon_{i}$ is a vanishing term that converges to zero as $N\to\infty$. Substituting this into the error update equation
$e_{j}^{(i,k+1)}=\epsilon_{i}+\sum_{l\in\mathcal{N}_{i}^{(k)}}w_{il}^{(k)}(e_{j}^{(l,k)}-e_{j}^{(i,k)}).$
We now analyze the behavior of the error dynamics. Denote the global error vector at time $k$ as
$\mathbf{e}_{j}^{(k)}=[e_{j}^{(1,k)},e_{j}^{(2,k)},\dots,e_{j}^{(m,k)}]^{T}.$
The update rule for the global error vector becomes
$\mathbf{e}_{j}^{(k+1)}=W_{k}\mathbf{e}_{j}^{(k)}+\boldsymbol{\epsilon},$
where $W_{k}$ is the doubly stochastic weighting matrix and $\boldsymbol{\epsilon}=[\epsilon_{1},\epsilon_{2},\dots,\epsilon_{m}]^{T}$. Taking the expectation of the error dynamics and using the fact that $W_{k}$ is doubly stochastic (so $W_{k}\mathbf{1}=\mathbf{1}$ and $W_{k}\mathbf{e}_{j}^{*}=\mathbf{e}_{j}^{*}$, where $\mathbf{e}_{j}^{*}$ is the true eigenvalue vector), we obtain
$\mathbb{E}[\mathbf{e}_{j}^{(k+1)}]=\overline{W}_{k}\mathbb{E}[\mathbf{e}_{j}^{(k)}]+\mathbb{E}[\boldsymbol{\epsilon}],$
where $\overline{W}_{k}=\mathbb{E}[W_{k}]$. Since $\epsilon_{i}\to 0$ as $N\to\infty$, we have $\mathbb{E}[\boldsymbol{\epsilon}]\to 0$. Now, observe that because $\overline{W}_{k}$ is doubly stochastic and the communication graph is connected, the matrix $\overline{W}_{k}$ has a spectral gap (i.e., its second-largest eigenvalue is less than 1), which implies that the errors $\mathbb{E}[\mathbf{e}_{j}^{(k)}]$ decay geometrically over time. Thus, as $k\to\infty$, the error terms vanish
$\lim_{k\to\infty}\mathbb{E}[\mathbf{e}_{j}^{(k)}]=0,$
which implies that
$\lim_{k\to\infty}\lambda_{j}^{(i,k)}=\lambda_{j}^{*},\quad\forall i.$
Hence, the consensus value converges to the true eigenvalue $\lambda_{j}^{*}$ as $k\to\infty$, completing the proof. ∎
The convergence rate depends on the spectral gap of the weight matrix $\mathbf{W}$; a larger spectral gap implies faster convergence. Therefore, optimizing $\alpha_{il}$ to maximize this gap can improve convergence speed.